OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

find out slope first!

OpenStudy (aroub):

1+3/2+1 =4/3 y=mx+b (2,1) 1=4/3(2)+b , 1=8/2+b, -8/2+2/2 =b -6/2=b, -3=b so it becomes y=4/3x-3